OpenStudy (anonymous):

Henry is putting a new baseboard around his room. He used the formula p=2(l+w) to find the perimeter. The perimeter is 72.5'. He remembers the width is 16.5'. Show 2 different ways to find the length of the other wall.

OpenStudy (anonymous):

Somebody Please Help!!!

OpenStudy (anonymous):

72.5=2*(l+16.5) 72.5/2=l+16.5 l=72.5/2 - 16.5

OpenStudy (anonymous):

i wouldnt know of a second way

OpenStudy (anonymous):

Thank You Very Much!!!